Drop a 4451 lb flat plate (area = 15444.5 sq inches) which stays flat while falling from 1 mile high. How much HP needed for a car to drive past impact point starting at zero and 1 mile from impact point? Known factors: Coefficent of drag C(d)= 1.28 , Air Density (no wind) p = .002377slugs per ft cubed. Gravity constant =g=32.2ft/sec squared, Horsepower needed to accelerate is AVERAGE - not peak
Answer #1
  • Gravitational force on the plate = mg.

    air drag force = 1/2 \rho v2Cd A.

  • area of the car A = 15444.5 /144 = 107.25 sq/ft

    Cd = 1.28
  • \rho = 0.002377 slug/cft = 0.002377*32.2 lb/cft
  • drag force Fd = 1/2 \rho v2Cd A. = 5.25v2

The plate is subjected to gravity and sir drag

its motion

ma = mg - Fd

dv/dt = g- 5.25v2

dv/( 1.87-v2) = 5.25 dt

integrating both sides we get :

use identity u”? – 70 – zB 0-2 T rp

V – 1.37 V +1.37 = -14.190

v(t) = 1371 +e-14.194 1-e14.19

  • When the plate reached terminal velocity both force on the car will be equal

    mg = 1/2 \rho v2Cd A.

    terminal vel . vt2 = (2mg/(\rhoCd A.) )

    = 2*4451/ (0.002377*1.28*107.25) ) = 27280

    vt = 165.17 ft/s

  • put v(t) = 165.17 in the above expression we get t = 0.001s

    it will reach terminal velocity at 0.001 s ,can be ignored

    time of fall t = 5280 /165.17 = 31.97 s

    You have to drive 1 mile (5280 ft) in less than 31.97 s to drive past before it falls to ground , starting with 0 vel

  • take it 31 s on safer side

    speed required = 5280/31 = 170.3 ft /s - average velocity

    we take mass of the car mass as 4451 lbs.

    KE of the car = 1/2 mv2 = 1/2 4451*(170.3)2 = 64,561,269 ft.lb

    you have to drive in 31 s

    1 HP = 550 ft.lb/s

    HP required to drive =64,561,269/550*31 = 3787 HP